# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ 4 Uniform Convergence Theorems for Net Riemann and $S^{*}$-partition Integrals
    Obtaining uniform convergence theorems for the Net Riemann integral, and specially the $S^{*}$(-partition) integral allows us to compare them to [30], [32] and to other order based sequential/net-based integrals in the literature (see the next section for a possible list of such integrals) and;
@@
    Based on the comparisons contained in item (i), we may obtain monotone convergence theorems and other limit type properties for these integrals, as well as initial steps for a general classification of such integration procedures.
To obtain these convergence theorems, we would like to make as few assumptions as possible on the integrability of the sequence of functions and (order based) limits involved. For that reason, we will adopt a generalization of the uniform convergence theorem of [48] to the Henstock-Kurzweil integral (with respect to the Lebesgue measure) which, compared to [10] - which extends the partially ordered operator-valued framework of [58] - posits less strong integrability conditions on the limit of the approximating functions. Again, we focus mainly on the Net Riemann integral, the modification for the Fleischer case being immediate.
We now fix some conventions for this section:
@@
    The existence of any integral with values on $Z$ will be understood to take place in the Boccuto-Candeloro net convergence structure given by $(D)$-convergence. All other types of limit will be specified.
@@
    The series appearing in the $S^{*}$-partition below will be understood to be taken in order convergence when the issue of existence arises. That will cause no problems with the requirement of $(D)$-convergence in $(i)$, as these will appear only when the functions and measures involved are non-negative and $Z$ weakly $\sigma$-distributive and Dedekind complete: this gives a increasing sequence (or net, in the $\mathcal{F}(\mathbb{N})$ case) of (unordered) sums which, by Propositions 2.5, 3.8 and 3.9 gives a well defined (and equal) limit in $(D)$-convergence as well. Therefore, we have compatibility with Definition 3.15 considering (i), where $\eta_{3}$ is $(D)$-convergence.
We begin by the following useful concept of [48]:
Definition 4.1 (Uniform Convergence in Order).
We say that a sequence of functions $f_{n}:\Omega\rightarrow X$ converges order uniformly to $f:\Omega\rightarrow X$ is there exists a non-negative and non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ such that $u_{n}\downarrow 0$ and:
$|f_{n}(x)-f(x)|\leq u_{n},$
for every $x\in\Omega$.
One important comment is that we may modify the definition above to make an "tail based" convergence concept (i.e, making the inequality valid for a "sufficient large nn", and not every $n$) based on $p_{n}$, but this modification will not cause any changes in the arguments below, as we can always make "tail convergence" modifications of these (that would not be the case if we consider uniform convergence based on nets of functions).
We also have the following Cauchy definition from [48]:
Definition 4.2 (Cauchy Sequence Uniformly in Order).
We say that a sequence of functions $f_{n}:\Omega\rightarrow X$ is uniformly Cauchy in order, or only uniformly Cauchy, if is there exists a non-negative and non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ such that $u_{n}\downarrow 0$ and for each $n\in\mathbb{N}$:
$|f_{i}(x)-f_{j}(x)|\leq u_{n},$
for every $x\in\Omega$ and $i,j\geq n$.
It is straightforward to check that every uniformly order convergent sequence of functions if uniformly Cauchy.
For the main convergence theorem, we need the following analogue of Proposition 3.2 of [48]:
@@
The proof is the same as in [48], but adapted to our abstract framework. Therefore, to present the modifications and for completeness, we give the full description.
@@
Now, as $\{f_{n}\}_{n\in\mathbb{N}}$ is uniformly Cauchy, there exists a non-increasing sequence $\{b_{n}\}_{n\in\mathbb{N}}$ consisting of elements of $X$ such that $u_{n}\downarrow 0$ and, for each $n\in\mathbb{N}$,
$|f_{i}(x)-f_{j}(x)|\leq u_{n},\ \forall x\in\Omega,\ i,j\geq k.$
@@
$f_{i}(x)\leq f_{j}(x)+u_{n},\forall x\in\Omega,$
and, therefore, as these functions are Net Riemann integrable, and the integral in this case is an isotone functional for the given bilinear product, we have that, integrating both sides of this inequality:
$\int_{\Omega}f_{i}d\mu\leq\int_{\Omega}f_{j}d\mu+u_{n}\mu(\Omega).$
In particular, from that it follows that:
$\int_{\Omega}f_{1}d\mu-u_{1}\mu(\Omega)\leq\int_{\Omega}f_{k}d\mu\leq\int_{\Omega}f_{1}d\mu+u_{1}\mu(\Omega),$ (1)
@@
The main result of this section, which uses the last result, is the analogue of Theorem 3.3 of [48], but modified for the purposes of the Net Riemann integral, which we now state:
@@
Again, in the proof below, we follow mainly [48], with mild modifications for the net Riemann type result. In particular, we will assume that $\Gamma$ is upwards directed, the downwards directed case being a straightforward modification.
@@
As $f_{n}$ converges to $f$ order uniformly, there exists a non-increasing sequence $\{u_{n}\}_{k\in\mathbb{N}}$ in $X$ such that $u_{n}\downarrow 0$ and:
$|f_{k}(x)-f(x)|\leq u_{n},\ \forall x\in\Omega\ \text{and}\ k\in\mathbb{N},$ (2)
@@
$\leq|\sum_{i\in I_{\gamma}}(f(\tau_{\gamma_{i}})-f_{n}(\tau_{\gamma_{i}}))\mu(\sigma_{\gamma_{i}})|+\bigvee_{i=1}^{\infty}a_{i\phi(i)}+w_{n}.$
Continuing from this last inequality we get that it is
$\leq\sum_{\alpha\in P}|f(\tau_{\gamma_{i}})-f_{n}(\tau_{\gamma_{i}})|\mu(\sigma_{\gamma_{i}})+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}$
@@
In fact, given $\varphi\in\mathbb{N}^{\mathbb{N}}$, let $p=\min_{i\in\mathbb{N}}\varphi(i+1)$ and $\psi\in\mathbb{N}^{\mathbb{N}}$ the function defined by $\psi(j)=\varphi(j+p+1)$, $\forall j\in\mathbb{N}$. By (5), for $\gamma\geq\gamma_{p}$, $\gamma_{p}=\gamma_{p}(\phi)$, all in $\Gamma$, we have that:
@@
Also, we have:
@@
4.1 Null sets and a Uniform Convergence Theorem for the S*-partition Integral
For this subsection, we will now present an uniform convergence for the $S^{*}$-partition integral outside of (to be defined) nulls sets of $\mu$. This will be used in an application bellow about the comparison of Riemann with Lebesgue type net based integrals In general, this type of result will not be true for the general Net Riemann integral which contains, in its special cases, integrals which behaves quite badly with respect to null sets integration - see, for exemple, the Henstock-Kurzweil case in [11]. Nevertheless, we also present such a theorem for the specific case of the Sion integral.
@@ 4.1 Null sets and a Uniform Convergence Theorem for the S*-partition Integral
Given a set function $\mu:\mathcal{H}\rightarrow Y$, we say that a set $A\in\mathcal{H}$ is $\mu$-null if $\mu(N)=0$.
@@
Let $\mu:\mathcal{H}\rightarrow Y^{+}$ be a positive additive set function. If $B\in\mathcal{H}$ is $\mu$-null and $\subseteq B, AA\in\mathcal{H}$, then $B$ is $\mu$-null
@@
In fact, we have that, by the non-negativity of $\mu$,
@@
Let $\mu:\mathcal{H}\rightarrow Y^{+}$, $N$ a $\mu-null$ set and $f:\Omega\rightarrow X$ a function. Then, $f$ is $S^{*}$-partition integrable on $N$ and:
@@
$|\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha)-0|\leq\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)},$
@@
Let $\mu:\mathcal{H}\rightarrow Y^{+}$, $N$ a $\mu-null$ set and $f:\Omega\rightarrow X$ a $S^{*}$-partition integrable function on $\Omega$. Then,
@@
$|f_{n}(x)-f(x)|\leq u_{n},\ \forall x\in\Omega\setminus N\ \text{and}\ n\in\mathbb{N}.$
@@
$|\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})|\leq\sum_{j\in J,\alpha\in P}|f(\delta_{\alpha_{j}})-f_{n}(\delta_{\alpha_{j}})|\mu(\alpha_{j})+\sum_{j\in J,\alpha\in P}f_{n}(\delta_{\alpha_{j}})\mu(\alpha_{j}),$
@@
$\sum_{j\in J,\alpha\in P}|f(\delta_{\alpha_{j}})-f_{n}(\delta_{\alpha_{j}})|\mu(\alpha_{j})\leq u_{n}\mu(\Omega\setminus N),$
@@
$\leq|\sum_{\alpha\in P}(f(\delta_{\alpha})-f_{n}(\delta_{\alpha}))\mu(\alpha)|+\bigvee_{i=1}^{\infty}a_{i\phi(i)}+w_{n}.$
Now, for each $P$ as above, the net $\{\sum_{j\in J,\alpha\in P}|f_{n}(\delta_{\alpha_{j}})-f(\delta_{\alpha_{j}})|\mu(\alpha_{j})\}_{J\in\mathcal{F}(\mathbb{N})}$ is order and $(D)$- convergent to its supremum, as it is increasing and bounded above by $u_{n}\mu(\Omega\setminus N)$ for each $n$ and $X$ is Dedekind complete. Therefore,
$\sum_{j\in J,\alpha\in P}|f_{n}(\delta_{\alpha_{j}})-f(\delta_{\alpha_{j}})|\mu(\alpha_{j})\leq\sum_{\alpha\in P}|f_{n}(\delta_{\alpha})-f(\delta_{\alpha})|\mu(\alpha_{j}),$
for each $J\in\mathcal{F}(\mathbb{N})$. Hence, by Proposition 1.20 of [44]:
$|\sum_{\alpha\in P}f_{n}(\delta_{\alpha})-f(\delta_{\alpha})\mu(\alpha_{j})|\leq\sum_{\alpha\in P}|f_{n}(\delta_{\alpha})-f(\delta_{\alpha})|\mu(\alpha_{j}),$
Therefore, from the last main inequality:
$\leq\sum_{\alpha\in P}|f_{n}(\delta_{\alpha})-f(\delta_{\alpha})|\mu(\alpha)+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}$
@@
Also, we have:
@@
Consider now $P_{0}^{{}^{\prime}}=P_{0}\cup\{N\}$, which is an element of $\mathcal{P}(\Omega)$ by construction. Take now any $P^{{}^{\prime}}\geq P_{0}{{}^{\prime}}$ in $\mathcal{P}(\Omega)$. Define the following in relation to $P^{{}^{\prime}}$:

### v2

@@ Abstract
    The class of sets $\mathcal{H}$ will be a $\sigma$-algebra. Nevertheless, many results are valid supposing that this class is only a $\delta$-ring.
@@
    The existence of any integral with values on $Z$, when it is a Riesz space, will be understood to take place in the net convergence structure given by $(D)$-convergence, unless stated to the contrary. All other types of limits will be specified.
  3. (iii)
    For $Z$ a Riesz space, the series appearing in the $S^{*}$-partition below will be understood to be taken in order convergence when the issue of existence arises. That will cause no problems with the requirement of $(D)$-convergence in $(i)$, as these will appear only when the functions and measures involved are non-negative and $Z$ weakly $\sigma$-distributive and Dedekind complete: this gives a increasing sequence (or net, in the $\mathcal{F}(\mathbb{N})$ case) of (unordered) sums which, by Propositions 1.5, 3.4 and 3.5 gives a well defined (and equal) limit in $(D)$-convergence as well. Therefore, we have compatibility with Definition 3.6 considering (i), where $\eta_{3}$ is $(D)$-convergence.
  4. (iv)
    Wherever we use the net Riemann integral, we assume that the Szabó-Száz-Fleischer net is composed of sets such that:
    $\sum_{i\in I_{\gamma}}\mu(\sigma_{\gamma_{i}})=\mu(\bigcup_{i\in I_{\gamma}}\sigma_{\gamma_{i}}),$
    for $\mu$ finitely additive, which is not the same as asking that the sequence of sets is disjoint, as it can have overlapping null sets - that is $A,B\in\mathcal{H}$ are not disjoint, but $A\cap B\in\mathcal{H}$ is such that $\mu(A\cap B)=0$, which does not alter any of the problems below.
We begin by the following useful concept of Monteiro and Fernandez (2012) for the order convergence theorems:
Definition 4.1 (Uniform Convergence in Order).
We say that a sequence of functions $f_{n}:\Omega\rightarrow X$ converges order uniformly to $f:\Omega\rightarrow X$ is there exists a non-negative and non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ in $X$ such that $u_{n}\downarrow 0$ and:
$|f_{n}(\omega)-f(\omega)|\leq u_{n},$
for every $\omega\in\Omega$.
One important comment is that we may modify the definition above to make a "tail based" convergence concept (i.e, making the inequality valid for a "sufficient large nn", and not every $n$) based on $p_{n}$, but this modification will not cause any changes in the arguments below, as we can always make "tail convergence" modifications of these (that would not be the case if we consider uniform convergence based on nets of functions).
We also have the following Cauchy definition from Monteiro and Fernandez (2012):
Definition 4.2 (Cauchy Sequence Uniformly in Order).
We say that a sequence of functions $f_{n}:\Omega\rightarrow X$ is uniformly Cauchy in order, or only uniformly Cauchy, if is there exists a non-negative and non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ in $X$ such that $u_{n}\downarrow 0$ and for each $n\in\mathbb{N}$:
$|f_{i}(\omega)-f_{j}(\omega)|\leq u_{n},$
for every $\omega\in\Omega$ and $i,j\geq n$.
It is straightforward to check that every uniformly order convergent sequence of functions is uniformly Cauchy.
We enounce the following direct consequence of Definition 2.4 and Proposition 1.1, which will be useful next.
Proposition 4.1.
Suppose $(Z,\eta_{3})$ is a Dedekind Complete Riesz space with $\eta_{3}$ being given by order convergence. Then the net Riemann integral of a function $f:\Omega\rightarrow X$ with respect to a set function $\mu:\mathcal{H}\rightarrow Y$ exists with respect to $\eta_{3}$ if and only if:
@@
    The net $\{S_{\gamma}(f,\mu)\}_{\gamma\in\Gamma}$ is bounded in $Z$ and,
@@
For the main convergence theorem, we need the following analogue of Proposition 3.2 of Monteiro and Fernandez (2012):
@@
The proof is the same as in Monteiro and Fernandez (2012), but adapted to our abstract framework. Therefore, to present the modifications and for completeness, we give the full description.
@@
Now, as $\{f_{n}\}_{n\in\mathbb{N}}$ is uniformly Cauchy, there exists a non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ consisting of elements of $X$ such that $u_{n}\downarrow 0$ and, for each $n\in\mathbb{N}$,
$|f_{i}(\omega)-f_{j}(\omega)|\leq u_{n},\ \forall\omega\in\Omega,\ i,j\geq k.$
@@
$f_{i}(\omega)\leq f_{j}(\omega)+u_{n},\forall\omega\in\Omega,$
and, therefore, as these functions are Net Riemann integrable, and the integral in this case is an isotone functional for the given bilinear product, we have that, integrating both sides of this inequality and using the regularity of the integrator:
$\int_{\Omega}f_{i}d\mu\leq\int_{\Omega}f_{j}d\mu+u_{n}\mu(\Omega).$ (3)
In particular, from that it follows that, for $j,n=1$:
$\int_{\Omega}f_{1}d\mu-u_{1}\mu(\Omega)\leq\int_{\Omega}f_{i}d\mu\leq\int_{\Omega}f_{1}d\mu+u_{1}\mu(\Omega),$ (4)
@@
The main result of this section, which uses the last result, is the analogue of Theorem 3.3 of Monteiro and Fernandez (2012), but modified for the purposes of the Net Riemann integral, which we now state:
@@
In the proof below, we follow mainly Monteiro and Fernandez (2012) and make some modifications inspired by the analogous result from Riečan and Vrábelová (1998). In particular, we will assume that $\Gamma$ is upwards directed, the downwards directed case being a straightforward modification.
@@
As $f_{n}$ converges to $f$ order uniformly, there exists a non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ in $X$ such that $u_{n}\downarrow 0$ and:
$|f_{k}(\omega)-f(\omega)|\leq u_{n},\ \forall\omega\in\Omega\ \text{and}\ k\in\mathbb{N},$ (5)
@@
$\leq|\sum_{i\in I_{\gamma}}(f(\tau_{\gamma_{i}})-f_{n}(\tau_{\gamma_{i}}))\mu(\sigma_{\gamma_{i}})|+\bigvee_{i=1}^{\infty}a_{i\psi(i)}+w_{n}.$
Continuing from this last inequality and, using 5 in the first term, we get that this last line is
@@
In fact, given $\varphi\in\mathbb{N}^{\mathbb{N}}$, let $p=\min_{i\in\mathbb{N}}\varphi(i+1)$ and $\psi\in\mathbb{N}^{\mathbb{N}}$ the function defined by $\psi(j)=\varphi(j+p+1)$, $\forall j\in\mathbb{N}$. By (8), for $\gamma\geq\gamma_{p}$, $\gamma_{p}=\gamma_{p}(\varphi)$, all in $\Gamma$, we have that:
@@
Also, we have, by the properties of the bilinear product:
@@
4.1 Null sets and a Uniform Convergence Theorem for the S*-partition Integral.
For this subsection, we will now present an uniform convergence for the $S^{*}$-partition integral outside of (to be defined) nulls sets of $\mu$. In general, this type of result will not be true for the general Net Riemann integral which contains, in its special cases, integrals which behaves quite badly with respect to null sets integration - see, for example, the Henstock-Kurzweil case in Boccuto et al. (2012b). Nevertheless, we also present such a theorem for the specific case of the Sion integral.
@@
Given a set function $\mu:\mathcal{H}\rightarrow Y$, we say that a set $N\in\mathcal{H}$ is $\mu$-null if $\mu(N)=0$.
@@
Let $\mu:\mathcal{H}\rightarrow Y^{+}$ be a non-negative (finitely) additive set function. If $B\in\mathcal{H}$ is $\mu$-null and $\subseteq B, AA\in\mathcal{H}$, then $B$ is $\mu$-null
@@
In fact, we have that, by the non-negativity and additivity of $\mu$,
@@
Let $\mu:\mathcal{H}\rightarrow Y^{+}$ (finitely) additive, $N$ a $\mu-null$ set and $f:\Omega\rightarrow X$ a function. Then, $f$ is $S^{*}$-partition integrable on $N$ and:
@@
$|\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha)|\leq\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)},$
@@
Let $\mu:\mathcal{H}\rightarrow Y^{+}$ (finitely) additive, $N$ a $\mu-null$ set and $f:\Omega\rightarrow X$ a $S^{*}$-partition integrable function on $\Omega$. Then,
@@
$|f_{n}(\omega)-f(\omega)|\leq u_{n},\ \forall\omega\in\Omega\setminus N\ \text{and}\ n\in\mathbb{N}.$
@@
$|\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})|\leq\sum_{j\in J,\alpha\in P}|(f(\delta_{\alpha_{j}})-f_{n}(\delta_{\alpha_{j}}))\mu(\alpha_{j})|+\sum_{j\in J,\alpha\in P}f_{n}(\delta_{\alpha_{j}})\mu(\alpha_{j}),$
@@
$\sum_{j\in J,\alpha\in P}|(f(\delta_{\alpha_{j}})-f_{n}(\delta_{\alpha_{j}}))\mu(\alpha_{j})|\leq u_{n}\mu(\Omega\setminus N),$
@@
$\leq|\sum_{\alpha\in P}(f(\delta_{\alpha})-f_{n}(\delta_{\alpha}))\mu(\alpha)|+\bigvee_{i=1}^{\infty}a_{i\phi(i)}+w_{n}.$ (16)
Now, for each $P$ as above, the net $\{\sum_{j\in J,\alpha\in P}|(f_{n}(\delta_{\alpha_{j}})-f(\delta_{\alpha_{j}}))\mu(\alpha_{j})|\}_{J\in\mathcal{F}(\mathbb{N})}$ is order and $(D)$- convergent to its supremum, as it is increasing and bounded above by $u_{n}\mu(\Omega\setminus N)$ for each $n$ and $X$ is Dedekind complete. Therefore,
$\sum_{j\in J,\alpha\in P}|(f_{n}(\delta_{\alpha_{j}})-f(\delta_{\alpha_{j}}))\mu(\alpha_{j})|\leq\sum_{\alpha\in P}|(f_{n}(\delta_{\alpha})-f(\delta_{\alpha}))\mu(\alpha_{j})|,$
for each $J\in\mathcal{F}(\mathbb{N})$. Hence, by Proposition 1.20 of Popov and Randrianantoanina (2013) and using that $|\cdot|$ is order continuous:
$|\sum_{\alpha\in P}(f_{n}(\delta_{\alpha})-f(\delta_{\alpha}))\mu(\alpha_{j})|\leq\sum_{\alpha\in P}|(f_{n}(\delta_{\alpha})-f(\delta_{\alpha}))\mu(\alpha_{j})|,$
Therefore, from (16):
$\leq\sum_{\alpha\in P}|(f_{n}(\delta_{\alpha})-f(\delta_{\alpha}))\mu(\alpha)|+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}$
@@
Also, we have, by the properties of the bilinear product:
@@
Consider now $P_{0}^{{}^{\prime}}=P_{0}\cup\{N\}$, which is an element of $\mathcal{P}(\Omega)$ by construction. Take now any $P^{{}^{\prime}}\geq P_{0}{{}^{\prime}}$ in $\mathcal{P}(\Omega)$. Define the following class of sets in relation to $P^{{}^{\prime}}$: